OpenStudy (anonymous):

Which is the equation of the line passing through (0, 8) and (4, 0)?

OpenStudy (anonymous):

wheres the equations

OpenStudy (anonymous):

y = -2x + 8 y = -2x + 4 y = -1/2x +8 y = -1/2x + 4

OpenStudy (anonymous):

y=-2x+8

OpenStudy (anonymous):

u would plug in 8=-2(0)+8 for both do the same

OpenStudy (anonymous):

@kevindragonlord